Suppose you walk 14.0 m straight west and then 22.0 m straight south. What is the compass direction of a line connecting your starting point to your final position? Give answer as an angle in degrees counterclockwise from the east axis.
Answer #1

14 Tan 2As per above figure, Tan A = 22/14

So, angle A = 52.5 deg

so angle wtih east axis = 180 + 52.5 = 232.5 deg Ans